OpenStudy (anonymous):

I am working an equation using quadratic formula to solve. i got a weird answer. I just need to know if I'm on the right track. The problem is 10xsquared + 2x -5 =0

hartnn (hartnn):

so what did u take the values of a,b,c as ?

OpenStudy (anonymous):

a=10, b=2, c=-5

hartnn (hartnn):

thats correct.. are u stuck at some step ? where ?

OpenStudy (anonymous):

At the end, I got 2 plus or minus the square root of 204 divided by 20. All my other square roots had worked out to be perfect squares, so I just was wondering am I doing something wrong.

OpenStudy (anonymous):

Negative 2

hartnn (hartnn):

\(\huge\frac{-2\pm \sqrt{204}}{20}\) if thats what u got then its correct

hartnn (hartnn):

but can be simplified!

OpenStudy (anonymous):

Woo hoo. O.k. Thanks. Do I have to simplifly 2 and 20 to 1 and 10?

hartnn (hartnn):

yes, like this : \(\huge\frac{-2\pm \sqrt{204}}{20}=\huge\frac{-1\pm \sqrt{51}}{10}\)

hartnn (hartnn):

did u get that^^ ?

OpenStudy (anonymous):

Yes. Thanks so much. Would you care to help me on one more. I will tell you the answer I got. I just need a good grade on this. I"m not doing well in math.

hartnn (hartnn):

sure, i will help :) just post it as another question,and close this one.